# Convergent transformations of visual representation in brains and models

### v1

@@ Abstract
A fundamental question in cognitive neuroscience is what shapes visual perception: the external world’s structure or the brain’s internal architecture. Although some perceptual variability can be traced to individual differences, brain responses to naturalistic stimuli evoke similar activity patterns across individuals, suggesting a convergent representational principle. Here, we test if this stimulus-driven convergence follows a common trajectory across people and deep neural networks (DNNs) during its transformation from sensory to high-level internal representations. We introduce a unified framework that traces representational flow by combining inter-subject similarity with alignment to model hierarchies. Applying this framework to three independent fMRI datasets of visual scene perception, we reveal a cortex-wide network, conserved across individuals, organized into two pathways: a medial-ventral stream for scene structure and a lateral-dorsal stream tuned for social and biological content. This functional organization is captured by the hierarchies of vision DNNs but not language models, reinforcing the specificity of the visual-to-semantic transformation. These findings show a convergent computational solution for visual encoding in both human and artificial vision, driven by the structure of the external world.
@@ 1 Introduction
Visual perception arises from the brain’s ability to internalize statistical regularities in the environment, transforming raw sensory input into structured, meaningful experience. Yet whether perception primarily reflects these external patterns or is shaped by the intrinsic computational architecture of the brain remains a central issue in cognitive science. This longstanding debate traces from philosophical perspectives like Plato’s allegory of the cave and theories of direct realism 1, which posit that perception is a direct reflection of the external world, to modern constructivist theories, which frame perception not as a passive imprint but as an active, model-based inference process 2, 3. A stimulus-driven view predicts that individuals will form convergent neural representations aligned with environmental structure4. Conversely, if internal constraints dominate, representations may diverge, reflecting idiosyncratic predictive models 5, 6, 7. Understanding the balance between these external and internal constraints is therefore foundational to any theory of cortical representation.
Empirical evidence supports both sides of the debate. Substantial individual variability has been observed in low-level perceptual processes such as motion detection, contrast sensitivity, and color perception, linked to differences in cortical anatomy and function 8, 9, 10. In contrast, studies using naturalistic paradigms show that when different individuals process identical stimuli, they display similar activity patterns, indicating a shared representational organization across brains 11, 12, 13, 14. This suggests that despite anatomical and experiential variability, common transformation pathways reliably lead neural computation toward convergent representational outcomes.
The principle of stimulus-driven convergence extends beyond the human brain to deep neural networks (DNNs). A substantial body of research has established a hierarchical correspondence between DNN layers and the processing stages of the visual system—a relationship that holds across diverse sensory modalities15, 16, 17, 18, 19. This alignment makes DNNs particularly effective probes for how sensory input is transformed into abstract, conceptual representations20, 21, 22. By providing an explicit, layer-by-layer mapping from low-level features to high-level semantics, DNNs offer a powerful model system for isolating the transformations that occur along the cortical hierarchy. Thus, characterizing a brain region’s alignment across the DNN hierarchy provides a functional fingerprint of its abstraction level and enables a data-driven approach for tracing information flow through the cortex.
Here, we investigated whether the transformation of visual information into abstract meaning unfolds along a shared, systematic trajectory across individuals and whether this trajectory is reflected in the representational dynamics of DNNs. To address this, we introduce a unified, data-driven framework that combines inter-subject alignment to identify cortical areas and networks with shared geometry, model-based alignment to map the hierarchy, and decomposition techniques to interpret the stimuli features driving the correspondence (see Fig. 1). We applied this framework to the massive-scale Natural Scenes Dataset (NSD) 23 and further validated the generalization of our findings in two independent fMRI datasets–BOLD5000 24 and THINGS-fMRI 25–with different image-viewing tasks.
Our analysis applied to scene perception revealed a cortex-wide representational network that is shared across individuals and organized into two principal streams with distinct functional roles. One stream follows a medial-ventral trajectory dedicated to encoding visually grounded scene structure, aligning with the classical ventral pathway 26, 27, 28. The other traces a lateral-dorsal route, selectively tuned to biological and social content, offering empirical support and a data-driven refinement for recent proposals of a ‘third visual pathway’ for social perception 29, 30. This convergence across brains and artificial systems provides a principled framework for mapping how stimulus-driven information propagates through cortical systems, revealing the functional architecture that supports the transformation of sensory information into abstract representations.
@@
Figure 1: A unified framework for tracing representational pathways. (A) Feature Extraction. For each image stimulus, we extracted equivalent representations from brain activity and deep neural networks. Single-trial fMRI responses were aggregated within HCP-MMP cortical parcels 31 to create vector representations of the brain’s response. Concurrently, layer-wise activations were extracted from pre-trained vision and language models to obtain representations across the full model hierarchies. Both brain and model vectors were used to compute Representational Dissimilarity Matrices (RDMs) based on their pairwise correlations. (B) Representational Alignment. Representational Similarity Analysis (RSA) 32 was used to compare RDMs, quantifying the alignment between brain parcels and model layers. (C) Representational Connectivity. Inter-parcel RSA was used to construct a cortical network based on shared representational geometry. Analyzing this network revealed the hierarchical flow of information and identified key representational hubs. (D) Shared Dimensions. Within these hubs, Kernel Multi-view Canonical Correlation Analysis (KMCCA) 33 was used to isolate a subspace of representational dimensions shared across all participants. This final step links the data-driven representational axes to specific properties of the scene content, allowing for a functional interpretation of the information driving the alignment across the network.
@@ 2 Results
2.1 Inter-subject convergence in stimulus representation
A central prediction of stimulus-driven perception is that different individuals processing the same stimuli should converge on a common representational geometry 13, 32. For the natural image-viewing tasks studied here, this implies that stimulus structure should drive representational alignment across observers in relevant cortical areas–from early visual cortex through high-level visual regions, and potentially into frontal areas implicated in visually guided decisions 34, 35. While inter-subject measures in naturalistic fMRI are often interpreted as ceiling estimates of signal-to-noise 23, 36, 37, we instead treated high inter-subject RSA (IS-RSA) as direct evidence that a region encodes stimulus information in a format shared across individuals. Identifying this cortical map of shared representational geometry thus provides a foundation for characterizing the spatial organization of stimulus-driven representations and serves as a starting point for examining how visual information is transformed across distinct brain regions.
To evaluate this prediction, we focused our analysis on the Natural Scenes Dataset (NSD) 23. This dataset comprises 7T fMRI recordings in which eight participants viewed thousands of naturalistic scenes images in a memory task, each presented up to three times across sessions spanning a full year. For each stimulus, we formed a beta vector for each HCP-MMP parcel by concatenating the single-trial responses of its voxels 31. We then computed a parcel-wise Representational Dissimilarity Matrix (RDM) for each subject and used Representational Similarity Analysis (RSA) to quantify the geometry alignment of these neural patterns 32. To isolate the stable, perception-driven component of the response from potential confounds like session timing, we computed inter-subject alignment (IS-RSA) using a ‘shifted-repetition’ matching. This method compares RDMs between subjects using responses to the same images but from different trial repetitions (e.g., the first repetition for one subject vs. the second for another), ensuring alignment is driven by stimulus content alone (see Supplementary Fig. A2 for a schematic; main pipeline in Fig. 1). Statistical significance was then assessed against a stimulus-shuffled null distribution (10,000 repetitions) to identify parcels with a significant shared geometry (see Methods).
IS-RSA captured significant inter-subject alignment throughout the visual hierarchy (Figs. 2a–b; see Supplementary Fig. D1 for full parcel-level statistics). We observed an alignment gradient that peaked in early visual cortex and extended into high-level ventral occipito-temporal and dorsal occipito-parietal regions–the main areas forming the visual system hierarchy34, 38, 35. These regions are also linked to high-level functions like scene perception and place memory 27, 39. Weaker, yet significant, alignment was also found in prefrontal cortex (PFC) areas typically linked to visual attention and task-dependent semantic processing and anterior areas associated to 40, 41. This spatial pattern is consistent with prior analyses of the NSD dataset using signal-to-noise and decoding methods23, 42, 43. Although some hemispheric asymmetries in alignment strength were present (Fig. 2b), the representational profiles of symmetric parcels in the left and right hemispheres were highly correlated, suggesting that both hemispheres encode a similar geometric structure (Supplementary Fig. A1). These results support the hypothesis that scene statistics impose a shared representational organization from low-level visual areas to higher-order association cortex.
Three anatomically clustered groups of parcels showed a consistent alignment across participants (Fig. 2b). The first is early visual cortex (V1–V4), region associated with processing low-level visual features 44. The second is a ventral hub spanning key regions for processing scene context and spatial layout (parcels: VMV1-3, PHA1-3) 27, 45, 28. The third is a dorsal hub in the lateral occipitotemporal cortex (LOTC), encompassing the MT+ complex and the temporoparietal junction (parcels: MT, MST, FST, V4t, TPOJ2-3). These regions, which constitute an extension of the dorsal stream beyond classic spatial processing, are typically linked to motion integration, social cognition, and multimodal processing 12, 30, 46, 39. In the following sections, we use these three hubs—early visual, ventral, and LOTC—as reference points for our subsequent analyses.
@@ 2.1 Inter-subject convergence in stimulus representation
Figure 2: Cortical distribution of representational alignment. (A) Inter-subject alignment (RSA, Pearson $r$) computed per parcel and grouped by macro-anatomical clusters 47. Box-plots show the ten clusters with the highest mean alignment in the NSD sample ($N=8$; symmetric HCP-MMP atlas 31). Red lines mark the parcel-wise null distribution (mean ± s.d.; 10 000 label permutations). All displayed parcels exceed chance (two-tailed, FDR-corrected) at $p<0.001$, except EC and Pfm ($p<0.01$), 31pd ($p<0.05$) and 23d/31pv (n.s.). Peak clusters include early visual cortex (V1–V4), the ventral hub (VMV1–3, PHA1–3) and the LOTC hub (V4t, MT, MST, FST, TPOJ2–3). Full-atlas results appear in Supplementary Fig. D1. (B–D) Surface maps of inter-subject (B), vision-model (C) and language-model (D) alignment. RSA values are averaged across subjects and—where applicable—across models. (E–F) Parcel-wise relation between inter-subject alignment and model–brain alignment for vision (E) and language (F). Vision parcels follow a power-law fit ($R^{2}=0.94$, 95% bootstrap CI); language parcels cluster near zero except for those in the LOTC hub. (G) Mean alignment by modality within the three hubs. Paired $t$-tests ($t(7)$, two-tailed, FDR-corrected) show higher vision-model alignment in early-visual and ventral hubs, and higher language-model alignment in the LOTC hub (all $p<0.0001$).
@@
2.2 Model–Brain Alignment Mirrors Inter-Subject Shared Geometry
Our previous analysis revealed a shared representational geometry across individuals; however, the nature of the information underlying this alignment remains unspecified. To determine whether this shared geometry is driven by low-level visual features or by more abstract, conceptual content, we compared the neural data to two distinct classes of computational models. Deep vision models, which learn representations directly from pixel data, provide a template for a visually-grounded processing hierarchy48, 49, 50. In contrast, large language models, which learn from text tokens, provide a template for modality-independent semantic and conceptual relationships. By systematically comparing cortical activity to both model types, we can therefore functionally characterize brain regions. This approach allows us to determine whether their organizing principles are primarily visual or more abstract in nature.
We analyzed a diverse collection of vision and language models based on Transformer and Vision Transformer (ViT) architectures, spanning supervised, self-supervised, and contrastive training objectives (see Table 1). For each model, we computed the layer-wise RSA between its layer activations and every HCP-MMP parcel (Fig. 1; Methods). To obtain a single summary measure of model-parcel correspondence, we then identified the maximum RSA value across all layers. This max-RSA value estimates the peak alignment achieved between a model and a brain region, irrespective of the specific hierarchical stage at which it occurs. This approach yields a cortex-wide alignment map directly comparable to the IS-RSA map (Fig. 2c–d). Statistical significance was assessed via a two-tailed permutation test (10,000 permutations; see Supplementary Fig. D1 for full atlas results).
Vision-trained models closely reproduced the inter-subject map. Across HCP parcels, a parcel’s vision-model alignment score scaled tightly with its IS-RSA value, following a power-law relationship ($R^{2}=0.94$;Fig. 2e; see Supl. analysis A.6 for details). Spatially, the cortical alignment map (Fig. 2c) mirrored the inter-subject pattern (Fig. 2b), with the highest alignment in the Early Visual, Ventral, and LOTC hubs, and significant, albeit weaker, alignment throughout the remaining visual areas and prefrontal cortex. This tight correspondence suggests that the cortical parcels where stimulus geometry is shared across individuals are precisely those whose representational structure is captured by image-trained networks, indicating that these models and the human brain converge on a common, stimulus-driven coding scheme for natural scenes.
Language models trained on image captions displayed a distinct and more localized alignment pattern. The cortical alignment map showed no strong alignment in early visual or ventral areas. Specifically, the classical visual streams exhibited negative RSA scores, indicating dissimilar representational geometries that contrast with visually grounded organizations (Fig.2d; Supplementary Fig.D2). In contrast, robust positive alignment emerged in the LOTC hub and neighboring parcels (Fig.2f), indicating that alignment in these regions is not visually grounded. A direct comparison of modalities revealed that language-model alignment exceeded vision-model alignment within this LOTC hub (Fig.2g).
These results serve as a modality control, indicating that while alignment in early and ventral areas is tied to visually grounded features, the information shared in the LOTC hub is organized around a non-visual structure captured by language embeddings of image descriptions. Beyond this modality dissociation, we next aimed to situate cortical regions along the visual-to-semantic transformation trajectory. To this end, we examined their layer-wise alignment profiles, using the model hierarchy as a proxy for the brain’s own computational stages.

### v2

# Shared representations in brains and models reveal a two-route cortical organization during scene perception

@@ Abstract
The brain transforms visual inputs into high-dimensional cortical representations that support diverse cognitive and behavioral goals. Characterizing how this information is organized and routed across the human brain is essential for understanding how we process complex visual scenes. Here, we applied representational similarity analysis to 7T fMRI data collected during natural scene viewing. We quantified representational geometry shared across individuals and compared it to hierarchical features from vision and language neural networks. This analysis revealed two distinct processing routes: a ventromedial pathway specialized for scene layout and environmental context, and a lateral occipitotemporal pathway selective for animate content. Vision models aligned with shared structure in both routes, whereas language models corresponded primarily with the lateral pathway. These findings refine classical visual-stream models by characterizing scene perception as a distributed cortical network with separable representational routes for context and animate content.
@@ 1 Introduction
Sensory inputs arrive at the brain as physical signals, but perception depends on how those signals are converted into structured patterns of cortical activity 37, 20. In vision, this conversion spans multiple stages, from early visual cortex to downstream regions whose activity patterns encode objects, scenes, and higher-order interactions 20, 35, 27, 77. Characterizing how these representations are transformed across the cortex is central to understanding how the brain performs this computation: identifying which regions encode specific aspects of the sensory input and mapping the networks that route this information between them.
Naturalistic neuroimaging provides a way to approach this problem. When different individuals view the same stimuli, activity patterns show reliable synchrony across widespread regions 42, 5, 12. This inter-subject synchrony is interpreted as evidence that, despite individual variability, a substantial stimulus-locked component is encoded similarly across brains. This shared component can be used to localize regions that reliably encode stimulus-driven information across individuals 62, 29. Representational geometry formalizes this idea by describing each region in terms of pairwise similarities among stimulus-evoked responses, enabling comparisons across individuals and regions, and relating representational variance to stimulus properties that organize it 55, 43, 85.
Representational geometry also enables direct comparisons between brains and deep neural networks (DNNs). Across vision and language models, internal activations can resemble cortical representations when both systems process the same stimuli 79, 10, 82, reflecting shared organizing dimensions that group stimuli according to similar coarse principles 13. In the visual domain, vision models show graded correspondence with cortical organization, with deeper layers capturing increasingly complex shape- and object-based features 51, 99, 17, 52. Complementarily, language models capture semantic regularities that abstract away from visual details, with recent work identifying correspondence in higher-level ventral and lateral regions where perceptual information interfaces with conceptual knowledge 72, 94. Comparing model-derived representations to brain responses therefore provides a compact way to situate regions along an ordered computational hierarchy and to distinguish visually grounded features from more abstract semantic representations 16, 23.
Here, we investigated how information about visual scenes is encoded and routed across cortex by analyzing representational geometry shared among individuals and DNNs. We introduced a unified framework that links shared representational structure across individuals, its correspondence to vision and language models, and the network topology implied by representational connectivity, while using dimensionality decomposition to relate these patterns to coarse stimulus properties (Fig. 1). Integrating these complementary approaches, which are often used in isolation, allowed us to move beyond localizing regions with shared information to characterizing how these regions organize as a distributed network and identifying the specific stimulus features that drive this organization.
We applied this framework to the Natural Scenes Dataset (NSD) 2, in which participants performed a memory recognition task while viewing thousands of natural scenes spanning diverse environments, objects, and social interactions. We recovered a cortex-wide organization with two dissociable processing routes, each specialized to represent distinct types of stimulus information. The medial-ventral stream was specialized for scene layout and environmental context, while the lateral stream was selectively tuned to animate and social content. DNNs captured the shared geometry of both routes, though with clear modality-specific differences: vision models aligned with both pathways, whereas language models corresponded primarily with the lateral stream. These findings are consistent with recent refinements of classical visual-system accounts 35, 54, including proposals of a ventromedial route for scene context and a lateral “third visual pathway” for social perception 78, 70.We then compared these results using two independent datasets (BOLD5000 11 and THINGS-fMRI 46) and found that, although the same cortical areas consistently emerged in the alignment analysis, replicating the full two-stream organization required rich scene and social content.
@@
Figure 1: Overview of the analysis pipeline. (A) Feature extraction. For each image stimulus, we extracted corresponding representations from brain activity and deep neural networks. Single-trial fMRI responses were aggregated within cortical parcels to create vector representations of the brain’s response. Concurrently, layer-wise activations were extracted from pre-trained vision and language models to obtain representations across the full model hierarchies. Both brain and model vectors were used to compute representational dissimilarity matrices (RDMs). Example input image adapted from Wikimedia (Bengt Nyman, $C$ BY 2.0). (B) Representational alignment. Representational Similarity Analysis (RSA) was used in two ways: (i) inter-subject RSA (IS-RSA), correlating parcel-wise RDMs across participants to estimate shared representational geometry; and (ii) brain–model RSA, correlating parcel RDMs with model-layer RDMs to quantify brain–model and layer-wise alignment profiles. (C) Representational connectivity. IS-RSA between pairs of parcels was used to construct a cortical network based on how similarly regions encode the stimulus set. The directionality of information flow was inferred from the peak model-layer alignment established in (B). (D) Shared dimensions. Within the main hubs identified in (C), Kernel Multi-view Canonical Correlation Analysis (KMCCA) was used to decompose the shared geometry into latent dimensions common across participants and to relate these dimensions to coarse scene features that drive alignment in different parts of the network.
@@ 2 Results
2.1 Inter-subject shared geometry during scene viewing
To localize cortical regions whose response patterns reflect reliable stimulus information in a shared format across people, we used inter-subject representational similarity analysis (IS-RSA) 29, 55. We applied this approach to the Natural Scenes Dataset (NSD) 2, which provides high-resolution 7T fMRI responses from eight participants performing an inter-session long-term recognition memory task in which they detected repeated images while viewing thousands of natural scenes. Our goal was to measure the stimulus-driven representational geometry related to visual content.
To isolate this geometry, we parcellated the cortical surface into 180 regions per hemisphere using the Human Connectome Project multimodal parcellation (HCP-MMP) atlas 34 and extracted the provided single-trial responses for each image. For each subject and parcel, we constructed representational dissimilarity matrices (RDMs) from pairwise distances between multivoxel patterns. We then quantified inter-subject alignment by correlating parcel-wise RDMs across participants, using a shifted-repetition scheme to reduce mnemonic and session-related confounds (Supplementary Note 1). This yielded a parcel-wise map of shared representational geometry.
IS-RSA captured significant alignment across occipital cortex and extended into higher-order association cortex (Fig. 2a; Supplementary Figs. 1-2). Alignment was strongest in early visual cortex (V1-V4) and extended into ventral occipitotemporal and dorsal occipitoparietal regions, consistent with hierarchical accounts of visual organization 20, 54, 28. We also observed significant, albeit weaker, alignment in prefrontal regions, which may reflect shared contributions of visual attention and task-dependent semantic processing30, 95, 8. Within this map, three anatomically clustered sets of parcels showed especially strong and consistent alignment across participants (Fig. 2b).
The first cluster corresponds to the early visual cortex (V1-V4), consistent with its role in encoding low-level visual features 93. The second comprised a ventral hub in ventromedial temporal cortex (VMV1-3, PHA1-3), overlapping scene-selective regions such as the parahippocampal place area (PPA) and adjacent cortex implicated in encoding scene layout and place memory 27, 77, 66. The third comprised a lateral occipitotemporal cortex (LOTC) hub including the motion-selective MT+ complex and posterior temporoparietal junction (MT, MST, FST, V4t, TPOJ2-3), near regions responsive to bodies and biological motion that have been implicated in social perception 70, 95, 6. In the following sections, we use these three hubs (early visual, ventral, and LOTC) as reference points for subsequent analyses.
This spatial distribution is consistent with prior NSD work reporting variability in signal reliability and decodability across cortex 2, 87, 18, 22. Across parcels, higher IS-RSA values were associated with higher signal reliability (Supplementary Note 2), and a within-subject RSA analysis comparing different repetitions within each participant produced a similar cortical pattern (Supplementary Fig. 3). Although alignment strength showed some hemispheric asymmetries (Fig. 2b), representational profiles of left-right homologous parcels were highly correlated, consistent with a largely bilateral organization (Supplementary Note 3). Together, these results indicate that IS-RSA isolates a distributed set of regions whose scene-evoked representational geometry is shared across individuals, providing a benchmark for model comparisons described in the next sections.
@@ 2.1 Inter-subject shared geometry during scene viewing
Figure 2: Inter-subject and model–brain alignment across cortex. (A) Inter-subject representational alignment (IS-RSA; Pearson’s $r$ between parcel-wise RDMs across participants) for parcels drawn from the macro-anatomical clusters with the highest mean IS-RSA (color legend). Boxplots show the distribution across subjects in the NSD sample ($N=8$, symmetric HCP–MMP atlas). Red lines indicate the parcel-wise null distribution (mean $\pm$ s.d.; 10,000 permutations). Unless otherwise annotated in the plot, parcels are significant at *** ($p<0.001$; two-tailed, FDR-corrected); additional annotations indicate ** ($p<0.01$), * ($p<0.05$), or n.s. ($p\geq 0.05$). Peak alignment occurs in early visual cortex (V1–V4), a ventral hub (VMV1–3, PHA1–3), and a lateral occipitotemporal (LOTC) hub (V4t, MT, MST, FST, TPOJ2–3). Full-atlas results are shown in Supplementary Fig. 1. (B–D) Cortical surface maps of alignment, averaged across subjects (and across models within each modality): (B) inter-subject IS-RSA; (C) vision–model RSA (maximum across layers, averaged across vision models); (D) language–model RSA (maximum across layers, averaged across language models). (E–F) Parcel-wise relationship between IS-RSA and model–brain alignment for (E) vision models and (F) language models. Points are colored by macro-anatomical group (as in panel A); other parcels are shown in white. Vision parcels follow an approximate power-law fit ($R^{2}=0.94$, shaded band: 95% bootstrap CI), whereas language-model alignment clusters near zero or negative except for parcels in and around the LOTC hub. (G) Modality comparison within the three hubs. Boxplots show hub alignment for vision and language models in early visual cortex, the ventral hub, and the LOTC hub (averaged across models). Paired two-tailed $t$-tests ($t(7)$, FDR-corrected) indicate stronger vision-model alignment in early visual and ventral hubs, and stronger language-model alignment in the LOTC hub.
@@
2.2 Model–brain alignment reveals modality-specific shared geometry
Although IS-RSA localized parcels whose representational geometry was shared across observers, it did not by itself specify which stimulus information underlay that shared structure. To distinguish visually grounded from more abstract features, we compared cortical RDMs with RDMs from pretrained vision and language models. Vision models provide hierarchies learned directly from pixels 101, 56, 98, whereas large language models (LLMs) provide a reference for text-derived structure. By contrasting brain alignment with these two model classes, we characterized whether a parcel’s shared geometry was better captured by visually grounded features or by higher-level language representations.
We analyzed 37 pretrained models based on Vision Transformer (ViT) 24 and Transformer 89 architectures, spanning supervised, self-supervised, and contrastive objectives (Table 1). For each model, we extracted layer-wise activations for each NSD image (vision models) or its caption (language models), constructed layer-specific RDMs, and quantified model-brain alignment by correlating each layer RDM with each parcel RDM. For each model, parcel-wise alignment was summarized as the maximum RSA across layers. Fig. 2c-d show per-parcel maps averaged across subjects and, within each modality, across models (significance was assessed with a permutation-based null distribution).
Vision models produced an alignment map that closely matched the IS-RSA spatial pattern (Fig. 2c vs. Fig. 2b). Parcels with high IS-RSA also showed strong vision-model alignment, with peaks in the Early Visual, Ventral, and LOTC hubs and weaker but significant effects across additional visual and prefrontal parcels. This overlap indicates that image-trained feature spaces capture a substantial component of the stimulus-structured geometry that is shared across individuals during scene viewing, consistent with prior NSD results and ventral-stream model-brain alignment work 51, 99, 17, 94, 2, 87. Across parcels, IS-RSA and vision-model alignment were strongly related by approximate power-law scaling ($R^{2}=0.94$; Fig. 2e), consistent with both measures being largely driven by a common stimulus-locked component of scene organization (see Supplementary Note 4 for analysis of scaling and reliability).
In contrast, autoregressive LLMs trained on next-token prediction showed alignment concentrated in LOTC and adjacent temporal cortex, and more weakly in anterior areas of visual cortex (Fig.2d,f), regions implicated in biological motion, social perception, and multimodal processing 72, 6, 70, 95. Alignment in these areas is consistent with prior reports of language-model correspondence in lateral/anterior temporal cortex and medial frontal regions in naturalistic paradigms associated with semantic processing 72, 94, 22. Within the LOTC hub, language-model alignment exceeded vision-model alignment (Fig.2g), indicating that the shared geometry expressed in this hub is better captured by caption-derived structure than by purely visual model features.
Conversely, early and ventral visual regions showed near-zero or negative RSA with these LLM embeddings (Fig. 2d; Supplementary Fig. 2). Negative RSA indicates a systematic mismatch in geometry: stimulus features that separate responses in these parcels, mainly dominated by visual characteristics, are not emphasized by the language-model representations. This lack of alignment contrasts with studies finding ventral correspondence using language models optimized for sentence similarity or language-vision contrastive objectives 94, 22. Unlike the next-token prediction models used here, similarity-based models organize embeddings in their output layer by semantic content–grouping descriptions of similar scenes–which results in geometries that overlap more substantially with high-level visual representations.